# FRAbench and GenEval: Scaling Fine-Grained Aspect Evaluation across Tasks, Modalities

### v1

@@ Abstract
Evaluating the open-ended outputs of large language models (LLMs) has become a bottleneck as model capabilities, task diversity, and modality coverage rapidly expand. Existing “LLM-as-a-Judge” evaluators are typically narrow in a few tasks, aspects, or modalities, and easily suffer from low consistency. In this paper, we argue that explicit, fine-grained aspect specification is the key to both generalizability and objectivity in automated evaluation. To do so, we introduce a hierarchical aspect taxonomy spanning 112 aspects that unifies evaluation across four representative settings — Natural Language Generation, Image Understanding, Image Generation, and Interleaved Text-and-Image Generation. Building on this taxonomy, we create FRAbench, a benchmark comprising 60.4k pairwise samples with 325k aspect-level labels obtained from a combination of human and LLM annotations. FRAbench provides the first large-scale, multi-modal resource for training and meta-evaluating fine-grained LMM judges. Leveraging FRAbench, we develop GenEval, a fine-grained evaluator generalizable across tasks and modalities. Experiments show that GenEval (i) attains high agreement with GPT-4o and expert annotators, (ii) transfers robustly to unseen tasks and modalities, and (iii) reveals systematic weaknesses of current LMMs on evaluation [^1] Our code can be found at https://github.com/gsfsdv/FRABench-and-GenEval.
@@ 1 Introduction
As large language models (LLMs) have shown amazing abilities in human-like question answering and creative generation [1, 2], assessing the quality of their free-form outputs has become increasingly challenging. Automated evaluators, a.k.a. “LLM-as-a-Judge” paradigm [3], have therefore received much research attention, with the purpose of providing absolute or relative scores sometimes accompanied by detailed feedback [4]. While existing evaluators [5, 6] are typically tailored to limited tasks, aspects, or modalities, which sharply constrains their scope of application. Equally troubling, the credibility of these evaluators remains in doubt — empirical evidence shows that even human annotators suffer from a low agreement in evaluating the open-ended responses of LLMs as task complexity rises, and current automated evaluators inherit the same vulnerability.
In this paper, we target a fine-grained evaluator designed to generalize across different tasks and modalities (i.e., text and images). We highlight that aspect-level assessment is the core of such generalizability. On one hand, aspects like engagement, naturalness, or creativity are inherently interrelated [7], thus transferring similar semantics and evaluation standards for various tasks. On the other hand, explicit aspect specification can reduce subjectivity. Our observations indicate that the low agreement among human annotators is mainly caused by their implicitly varying weighting over aspects, rather than disagreement over the definition of any single aspect. For example, when judging story generation, some prioritize creativity while others emphasize logical consistency. By specifying aspects, one can substantially mitigate this variance source and obtain more objective evaluations.
@@
GenEval (ours) 112 Synthetic & Human
@@
To this end, we first build a fine-grained aspect tree by conducting a comprehensive survey on four types of (multi-modal) tasks: Natural Language Generation (NLG), Image Understanding (IU), Image Generation (IG), and Interleaved Text-and-Image Generation (ITIG). We then manually organize, extend, and redefine 112 aspects with hierarchical relations as a universal evaluation taxonomy—aspect tree. Note that although recent studies like Prometheus [12] and BIGGen-Bench [13] use auto-generated aspects for ultra-fine evaluation, we avoid this approach for two reasons. First, when the aspect becomes that narrowly tailored, its boundaries with direct evaluation become blurred, thereby reintroducing the very subjectivity we aim to suppress. Second, the universe of evaluation aspects is comparatively stable and, in practice, enumerable. Once the application scenario is specified, human or automated evaluators can simply traverse our aspect tree to select the relevant ones. This design dramatically lowers the cognitive load of defining evaluation standards and promotes consistency across tasks and modalities.
@@
In summary, our contributions are as follows: (1) We construct an evaluation aspect tree covering 112 aspects to ease fine-grained evaluation. (2) We develop FRABench, the first large-scale, multi-modal resource for training and meta-evaluating fine-grained LLM judges. (3) We create GenEval and conduct experiments to showcase its generalization capabilities, providing key insights.
@@ 2.2 Aspect Taxonomy Construction
In this section, we construct an aspect tree based on the aspects to establish a standardized evaluation taxonomy for multimodal evaluation tasks. Based on this universal taxonomy, both human and automated evaluators can easily traverse the aspect tree to customize evaluation criteria for specific tasks or scenarios, enabling fine-grained and objective assessments.
We first use the overall aspect as the root node of the aspect tree, and then divide the remaining aspects into two subtrees based on their task independence and task specificity:
  • Universal Aspects (UAs): These aspects exhibit task independence, as they primarily focus on assessing the quality of the model’s output. They are often modality-specific (for example, fluency for text or fidelity for image) and serve as fundamental quality aspects across tasks.
  • Task-specific Aspects (TAs): These aspects exhibit task dependence, as they primarily focus on task completion rather than output quality. They are typically closely tied to the task type. For example, engagingness for story generation or accuracy for mathematical reasoning.
For the hierarchical construction within the two subtrees, we first obtain existing tree structures from previous studies [14, 15, 16, 17, 18, 19]. We then classify these hierarchical structures, together with aspects lacking predefined hierarchies, into UAs and TAs as described above. Finally, for aspects without predefined hierarchies, we determine their positions within the tree according to their definitions—either assigning them to an appropriate location within an existing tree or establishing them as new root nodes. For example, the definition of alignment encompasses both object alignment and attribute alignment. This definition-driven approach is reasonable, as it ensures that the constructed hierarchy accurately reflects the semantic relationships among aspects. The final two subtrees are shown in Figure 2 and Figure 2.
Based on the aspect tree, the aspect-level evaluation process can be streamlined. Specifically, we first determine the correspondence between each UAs subtree (e.g, the subtree rooted at readability, represented by readability) and its respective modality: readability for text, image quality for image, image coherence for multi-images, and text-image relationship for text-with-image, while bias and toxicity apply across all modalities. Once the evaluation task is determined, we can identify the corresponding subtree based on the output modality to select appropriate aspects. Additionally, task-specific aspects can be traversed from the TAs according to the objectives of the task.
@@ 5.2 In-Domain Evaluation
The results demonstrate that GenEval consistently achieves higher alignment with both GPT-4o and human annotators compared to the baselines (see shaded area in radar chart). Specifically, excluding GPT-4o, GenEval surpasses all baselines on NLG, IU, and ITIG tasks within FRA-ID and FRA-ID-H, achieving overall average accuracies of 88.7% and 74.6%, respectively. On NLG, GenEval outperforms Themis, while our base model Qwen performs much worse, suggesting that with appropriate training, LMM-based evaluators can also effectively assess NLG tasks. Moreover, GenEval shows strong performance on the ITIG of FRA-ID-H, where its alignment with human annotators exceeds that of GPT-4o. This improvement can be attributed to GenEval’s generalization ability and cross-task transfer performance. For aspect-level comparison, GenEval outperforms GPT-4o on some aspects, with two illustrative examples shown in Figure 5a.
@@
Figure 6: Comparison on FRA-OOD and FRA-OOD-H. Black sectors show unseen TAs (for task and aspect generalization), while the rest show seen UAs for unseen tasks (for task generalization).
@@
Table 3: Average accuracy on FRA-OOD and FRA-OOD-H. Seen UAs for Unseen Tasks are used to assess task generalization, while Unseen TAs for Unseen Tasks evaluate aspect generalization.
@@ 5.3.1 Tasks Generalization Evaluation
We use FRA-OOD and FRA-OOD-H to evaluate GenEval’s generalization to tasks. The results are shown in the regions outside the black sectors in Figure 6. GenEval demonstrates strong alignment with both GPT-4o and human annotators on all tasks. In the “Seen UAs for Unseen Tasks” column of Table 3, GenEval achieves overall average accuracies of 85% and 83% on FRA-OOD and FRA-OOD-H, respectively, and outperforms both Themis and LLaVA-Critic on all tasks. Supporting the effectiveness of aspect-level evaluation for cross-task generalization. As illustrated in Figure 5a for aspect-level comparison, GenEval surpasses the baselines across most aspects, with only five aspects displayed here. We also provide several good cases in the Appendix G.
@@ 5.3.2 Aspects Generalization Evaluation
We use the unseen TAs in FRA-OOD and FRA-OOD-H to evaluate GenEval’s generalization to aspects. To support the validity of aspect generalization evaluation, we provide ROUGE_L [30] tests on unseen aspects in Appendix B.2. The results are shown in the black sectors of Figure 6, and in Table 3. In terms of overall coverage, GenEval outperforms both Themis and LLaVA-Critic, the state-of-the-art evaluators in their respective domains. It continues to demonstrate strong overall performance on the NLG, IU, and ITIG, with high consistency with both GPT-4o and human annotations, achieving average accuracies of 86.2% and 83.2%, respectively. For aspect-level comparison, as shown in Figure 5a, GenEval outperforms GPT-4o on unseen aspects such as suitability and explainability under NLG and IU in FRA-OOD-H. Additionally, for UAs generalization evaluation, since GenEval has already transferred relevant UAs from other tasks, it performs well on FRAUAs-OOD. Experimental results are provided in Appendix D.1. These results demonstrate that even without exposure to these unseen aspects during training, GenEval remains effective in assessment. This is largely due to the inherently interrelated nature of aspects, which enables the transfer of similar semantics and evaluation standards across different tasks.
@@ Appendix A Aspect Sources and Definitions
We provide the sources and definitions of aspects under two categories: UAs and TAs, as shown in Tables 13 to 26. For UAs, the “Target” column indicates the applicable output modality: “T’ for Text, “I” for Image, and “ITI” for Interleaved Text-with-Image. For TAs, the “Sub-Task” column specifies the corresponding task type, as defined in Figure 3. Additionally, underscores are used to benchmark extended TAs for ITIG.
@@ Appendix B Analysis of the FRABench
In this section, we provide detailed information and comprehensive analysis about the FRABench.
@@ B.3 Position Bias Analysis
We provide the counts of evaluation samples in the training set, FRA-ID, FRAUAs-OOD, and FRA-OOD, showing cases where Response 1 is better than, equal to, or worse than Response 2, as illustrated in Figures 10 to 12. As shown in Figure 10, the four tasks in the training set maintain a relatively consistent distribution. In Table 11, we also evaluate the position consistency of GenEval on the FRA-ID and FRA-OOD. We perform two inferences on the same sample, swapping the positions of the responses in the second inference, and calculate the average consistency, i.e., whether the results from the two inferences align. These experimental results demonstrate that GenEval achieves higher positional consistency compared to other baselines.
@@
Figure 9: Statistics of pairwise data in the training set.
Figure 10: Statistics of pairwise data in the FRAUAs-OOD.
Figure 11: Statistics of pairwise data in the FRA-ID.
Figure 12: Statistics of pairwise data in the FRA-OOD.
@@ B.4 Aspect Distribution Across Different Sub-tasks
Table 12: Aspects and sample counts in the Training Set, FRA-ID, FRAUAs-OOD, and FRA-OOD. Blue-marked aspects are human-annotated, while the rest are annotated by GPT-4o.
@@ C.3 Aspects Used in the benchmark
Since GenEval requires aspects for evaluation, we assign one to each benchmark. Specifically, we select suitable aspects from the TAs and UAs, and for unseen tasks, we define appropriate aspects based on task properties. Tables 27 to 29 show the aspects used by GenEval, with blue indicating unseen aspects. GPT-4o, Claude-3.5, Qwen2VL-7B, and Qwen2VL-72B also follow the same aspect assignments.
@@
Table 27: Aspects used by evaluators in the benchmark evaluations. (Part 1)
@@
Table 28: Aspects used by evaluators in the benchmark evaluations. (Part 2)
@@
Table 29: Aspects used by evaluators in the benchmark evaluations. (Part 3)
@@ D.1 Specific Experimental Results
We report the specific average accuracy of each evaluator for every aspect within each task in Figures 4 and 6. The alignment is measured against: (1) GPT-4o on the FRA-ID and FRA-OOD, and (2) human annotators on the FRA-ID-H and FRA-OOD-H. Detailed results are presented in Tables 30 to 41. The reported average accuracy is computed by aggregating the same aspect across different sub-tasks within the same task. We also include the comprehensive results on MLLM-as-a-Judge in Table 47. The experimental results of FRAUAs-OOD are presented in Figure 13 and Table 46.
@@
Figure 13: Comparison on FRAUAs-OOD to evaluate the generalization in UAs.
Method GPT-4o Alignment
@@
Table 46: Average accuracy on FRAUAs-OOD.
@@
D.2 Phased Training Experiment Results
During GenEval training, we compared unified training and phased training across tasks. For phased training, we group tasks with similar output modalities: NLG and IU as Text, IG as Image, and ITIG as Inter. Using the fine-tuning method from Appendix C.1, we train models in phases, swapping the order of these task types, resulting in six models. Experiments on the FRA-ID and FRA-ID-H (Figure 14 and Table 48) show that the model performs best on tasks from the last-trained category. In contrast, unified training provided more balanced performance across tasks.
Figure 14: Experimental results in the FRA-ID and FRA-ID-H using different types of training set orders for phased training.
Method GPT-4o Alignment Human Alignment

### v2

@@ Abstract
Evaluating the open-ended outputs of large language models (LLMs) has become a bottleneck as model capabilities, task diversity, and modality coverage rapidly expand. Existing “LLM-as-a-Judge” evaluators are typically narrow in a few tasks, aspects, or modalities, and easily suffer from low consistency. In this paper, we argue that explicit, fine-grained aspect specification is the key to both generalizability and objectivity in automated evaluation. To this end, we propose a hierarchical aspect taxonomy encompassing 112 distinct aspects that unifies evaluation across four representative settings — Natural Language Generation, Image Understanding, Image Generation, and Interleaved Text-and-Image Generation. Building upon this taxonomy, we create FRABench, a benchmark comprising 60.4k pairwise samples with 325k evaluation labels obtained from a combination of human and LLM annotations. FRABench provides the first large-scale, multi-modal resource for training and meta-evaluating fine-grained LMM judges. Leveraging FRABench, we develop GenEval, a fine-grained evaluator generalizable across tasks and modalities. Experiments show that GenEval (i) attains high agreement with GPT-4o and expert annotators, (ii) transfers robustly to unseen tasks and modalities, and (iii) reveals systematic weaknesses of current LMMs on evaluation [^1] Our code can be found at https://github.com/ALEX-nlp/FRABench-and-GenEval.
@@ 1 Introduction
As large language models (LLMs) have shown amazing abilities in human-like question answering and creative generation [1, 2], assessing the quality of their free-form outputs has become increasingly challenging. Automated evaluators, a.k.a. “LLM-as-a-Judge” paradigm [3], have therefore received much research attention, with the purpose of providing absolute or relative scores sometimes accompanied by detailed feedback [4]. While existing evaluators [5, 6, 7] are typically tailored to limited tasks, aspects, or modalities, which sharply constrains their scope of application. Equally troubling, the credibility of these evaluators remains in doubt — empirical evidence shows that even human annotators suffer from a low agreement in evaluating the open-ended responses of LLMs as task complexity rises, and current automated evaluators inherit the same vulnerability.
In this paper, we target a fine-grained evaluator designed to generalize across different tasks and modalities (i.e., text and image). We highlight that aspect-level assessment is the core of generalizability. Because aspects like engagement, naturalness, or creativity are inherently interrelated [8], thus transferring similar semantics and evaluation standards for various tasks. Additionally, explicit aspect specification can also reduce subjectivity. Our observations indicate that the low agreement among human annotators is mainly caused by their implicitly varying weighting over aspects, rather than disagreement over the definition of any single aspect. For example, when judging story generation, some prioritize creativity while others emphasize logical consistency. By specifying aspects, one can substantially mitigate this variance source and obtain more objective evaluations.
@@
GenEval 112 Synthetic & Human
@@
To this end, we first conduct a comprehensive survey on four types of (multi-modal) tasks: Natural Language Generation (NLG), Image Understanding (IU), Image Generation (IG), and Interleaved Text-and-Image Generation (ITIG) to identify key evaluation aspects. We then manually organize, extend, and redefine 112 distinct aspects with hierarchical relations as a universal evaluation taxonomy—aspect tree. Note that although recent studies like Prometheus [13] and BIGGen-Bench [14] use auto-generated aspects for ultra-fine evaluation, we avoid this approach for two reasons. First, when the aspect becomes that narrowly tailored, its boundaries with direct evaluation become blurred, thereby reintroducing the very subjectivity we aim to suppress. Second, the universe of evaluation aspects is comparatively stable and, in practice, enumerable. Once the application scenario is specified, human or automated evaluators can simply traverse our aspect tree to select the relevant ones. This design dramatically lowers the cognitive load of defining evaluation standards and promotes consistency across tasks and modalities.
@@
In summary, our contributions are as follows: (1) We construct an evaluation aspect tree covering 112 aspects to ease fine-grained evaluation. (2) We develop FRABench, the first large-scale, multi-modal resource for training and meta-evaluating fine-grained LMM judges. (3) We create GenEval and conduct experiments to showcase its generalization capabilities, providing key insights.
@@ 2.2 Aspect Taxonomy Construction
In this section, we construct an aspect tree based on the aspects to establish a standardized evaluation taxonomy for multimodal evaluation tasks. Based on this universal taxonomy, both human and automated evaluators can easily traverse the aspect tree to customize evaluation settings for specific tasks or scenarios, enabling fine-grained and objective assessments.
We first designate the overall aspect as the root node and construct two subtrees, then categorize the remaining aspects into corresponding subtrees based on task independence and specificity:
  • Universal Aspects (UAs): Aspects in this subtree exhibit task independence, as they primarily focus on assessing the quality of the model’s output. They are often modality-specific (for example, fluency for text or fidelity for image) and serve as fundamental quality aspects across tasks.
  • Task-specific Aspects (TAs): Aspects in this subtree exhibit task dependence, as they primarily focus on task completion rather than output quality. They are typically closely tied to the task type. For example, engagingness for story generation or accuracy for mathematical reasoning.
For the hierarchical construction within the two subtrees, we first adopt existing aspect trees from prior studies [15, 16, 17, 18, 19, 20]. For aspects lacking predefined hierarchies, we adopt an iterative methodology where, in accordance with the definitions, the aspect is either (1) integrated into an existing tree (for example, the definition of alignment encompasses both object alignment and attribute alignment.) or (2) created as the root node of a new tree. This definition-driven approach is reasonable, as it ensures that the constructed hierarchy accurately reflects the semantic relationships among aspects. Finally, each tree’s root nodes serve as child nodes of the corresponding two subtrees. The final two subtrees are shown in Figure 2 and Figure 2.
Based on the aspect tree, the aspect-level evaluation process can be streamlined. Specifically, we first determine the correspondence between each subtree (e.g, the subtree rooted at readability, represented by readability) and its respective modality: readability for text, image quality for image, image coherence for multi-images, and text-image relationship for text-with-image, while bias and toxicity apply across all modalities. Once the evaluation task is determined, we can identify the corresponding subtree based on the output modality to select appropriate UAs. Additionally, task-specific aspects can be traversed from the TAs according to the objectives of the task.
@@ 5.2 In-Domain Evaluation
The results demonstrate that GenEval consistently achieves higher alignment with both GPT-4o and human annotators compared to the baselines (see shaded area in radar chart). Specifically, excluding GPT-4o, GenEval surpasses all baselines on NLG, IU, and ITIG tasks within FRA-ID and FRA-ID-H, achieving overall average accuracies of 82.4% and 72.1%, respectively. On NLG, GenEval outperforms Themis, while our base model Qwen2Vl-7B performs much worse, suggesting that with appropriate training, LMM-based evaluators can also effectively assess NLG tasks. Moreover, GenEval shows strong performance on the ITIG of FRA-ID-H, where its alignment with human annotators exceeds that of GPT-4o. This improvement can be attributed to GenEval’s generalization ability and cross-task transfer performance. For aspect-level comparison, GenEval outperforms GPT-4o on some aspects, with two illustrative examples shown in Figure 5a.
@@
Figure 6: Comparison on FRA-OOD and FRA-OOD-H. Black sectors show unseen TAs (for task generalization), while the rest show seen UAs for unseen tasks (for task generalization).
@@
Table 3: Average accuracy on FRA-OOD and FRA-OOD-H. “Seen UAs for Unseen Tasks” are used to assess task generalization, while “Unseen TAs for Unseen Tasks” evaluate aspect generalization.
@@ 5.3.1 Tasks Generalization Evaluation
We use FRA-OOD and FRA-OOD-H to evaluate GenEval’s generalization to tasks. The results are shown in the regions outside the black sectors in Figure 6. GenEval demonstrates strong alignment with both GPT-4o and human annotators on all tasks. In the “Seen UAs for Unseen Tasks” column of Table 3, GenEval achieves overall average accuracies of 81% and 77.9% on FRA-OOD and FRA-OOD-H, respectively, and outperforms both Themis and LLaVA-Critic on all tasks. Supporting the effectiveness of aspect-level evaluation for cross-task generalization. As illustrated in Figure 5a for aspect-level comparison, GenEval surpasses the baselines across most aspects, with only five aspects displayed here. We also provide several good cases in the Appendix G.
@@ 5.3.2 Aspects Generalization Evaluation
We use the unseen TAs in FRA-OOD and FRA-OOD-H to evaluate GenEval’s generalization to aspects. To support the validity of aspect generalization evaluation, we provide ROUGE_L [31] tests on unseen aspects in Appendix B.2. The results are shown in the black sectors of Figure 6, and in Table 3. In terms of overall coverage, GenEval outperforms both Themis and LLaVA-Critic, the state-of-the-art evaluators in their respective domains. It continues to demonstrate strong overall performance on the NLG, IU, and ITIG, with high consistency with both GPT-4o and human annotations, achieving overall average accuracies of 80.3% and 78.9%, respectively. For aspect-level comparison, as shown in Figure 5a, GenEval outperforms GPT-4o on unseen aspects such as suitability and explainability under NLG and IU in FRA-OOD-H. Additionally, for UAs generalization evaluation, since GenEval has already transferred relevant UAs from other tasks, it performs well on FRAUAs-OOD. Experimental results are provided in Appendix D.1. These results demonstrate that even without exposure to these unseen aspects during training, GenEval remains effective in assessment. This is largely due to the inherently interrelated nature of aspects, which enables the transfer of similar semantics and evaluation standards across tasks and modalities.
@@ Appendix A Aspect Sources and Definitions
We provide the sources and definitions of aspects under two categories: UAs and TAs, as shown in Table 13 to 26. For UAs, the “Target” column indicates the applicable output modality: “T’ for Text, “I” for Image, and “ITI” for Interleaved Text-with-Image. For TAs, the “Sub-Task” column specifies the corresponding task type, as defined in Figure 3. Additionally, underscores are used to benchmark extended TAs for ITIG.
@@ Appendix B Analysis of the FRABench
In this section, we provide detailed information and comprehensive analysis about the FRABench. The dataset is currently available via an anonymous link: https://github.com/GenEval-1/GenEval.
@@ B.3 Position Bias Analysis
We provide the counts of evaluation labels in the training set, FRA-ID, FRAUAs-OOD, and FRA-OOD, showing cases where Response 1 is better than, equal to, or worse than Response 2, as illustrated in Figures 10 to 12. As shown in Figure 10, the four tasks in the training set maintain a relatively consistent distribution. In Table 11, we also evaluate the position consistency of GenEval on the FRA-ID and FRA-OOD. We perform two inferences on the same sample, swapping the positions of the responses in the second inference, and calculate the average consistency, i.e., whether the results from the two inferences align. These experimental results demonstrate that GenEval achieves higher positional consistency compared to other baselines.
@@
Figure 9: Statistics of pairwise evaluation labels in the training set.
Figure 10: Statistics of pairwise evaluation labels in the FRAUAs-OOD.
Figure 11: Statistics of pairwise evaluation labels in the FRA-ID.
Figure 12: Statistics of pairwise evaluation labels in the FRA-OOD.
@@ B.4 Aspect Distribution Across Different Sub-tasks
Table 12: Aspects and evaluation label counts in the Training Set, FRA-ID, FRAUAs-OOD, and FRA-OOD. Blue-marked aspects are human-annotated, while the rest are annotated by GPT-4o.
@@ C.3 Aspects Used in the benchmark
Since GenEval requires aspects for evaluation, we assign one to each benchmark. Specifically, we select suitable aspects from the TAs and UAs, and for unseen tasks, we define appropriate TAs based on task properties. Table 27 to 29 show the aspects used by GenEval, with blue indicating unseen aspects. GPT-4o, Claude-3.5, Qwen2VL-7B, and Qwen2VL-72B also follow the same aspect assignments.
@@
Table 27: Aspects used by evaluators in the public benchmark evaluations. (Part 1)
@@
Table 28: Aspects used by evaluators in the public benchmark evaluations. (Part 2)
@@
Table 29: Aspects used by evaluators in the public benchmark evaluations. (Part 3)
@@ D.1 Specific Experimental Results
We report the specific average accuracy of each evaluator for every aspect within each task in Figures 4 and 6. The alignment is measured against: (1) GPT-4o on the FRA-ID and FRA-OOD, and (2) human annotators on the FRA-ID-H and FRA-OOD-H. Detailed results are presented in Table 30 to 41. The reported average accuracy is computed by aggregating the same aspect across different sub-tasks within the same task. We also include the comprehensive results on MLLM-as-a-Judge in Table 50. The experimental results of FRAUAs-OOD are presented in Figure 13 and Table 46 to 49.
@@
Figure 13: Comparison on the FRAUAs-OOD to evaluate the generalization in UAs.
Method GPT-4o (FRAUAs-OOD)
@@
Table 46: Comparison on FRAUAs-OOD. We compute the average accuracy of UAs for each task.
Aspect Abbr GPT-4o Claude-3.5 Qwen2VL-72B Qwen2VL-7B Themis Ours
Clarity Cla 90.0 82.5 65.8 42.5 36.7 84.2
Coherence Coh 89.2 90.8 78.3 50.0 40.8 83.3
Fluency Flu 91.7 89.2 71.7 49.2 56.5 81.7
Grammaticality Gra 86.7 51.0 65.8 50.0 68.8 82.9
Harmfulness Har 92.4 81.0 90.9 65.3 80.6 93.9
Readability Rea 89.4 86.7 72.0 48.0 38.6 85.5
Sentence Complexity SenC 89.7 75.3 60.8 52.7 38.5 82.0
Spelling Accuracy SpeA 91.2 84.1 78.0 38.9 78.9 86.3
Structure Accuracy StrA 82.4 74.4 63.4 53.3 52.4 80.2
Tense Consistency TenC 96.2 61.5 80.8 64.0 73.1 80.8
Text Repetition TexR 90.0 78.5 52.0 38.2 41.5 77.5
Vocabulary Complexity VocC 81.7 72.6 57.5 47.6 44.7 75.5
Table 47: The specific experimental results of the NLG task in the FRAUAs-OOD are used to evaluate generalization in UAs.
Bias Bias 88.3 86.4 82.6 64.3 23.8 86.2
Clarity Cla 75.2 63.5 53.7 51.0 49.5 66.7
Coherence Coh 91.8 87.3 73.3 50.3 73.8 89.7
Ethnic Bias EthB 99.5 98.5 98.0 83.9 14.7 99.8
Fluency Flu 81.6 76.0 71.3 52.6 67.2 80.0
Grammaticality Gra 90.5 56.8 83.0 65.0 37.7 89.7
Readability Rea 88.8 84.0 70.6 50.7 72.6 85.2
Sentence Complexity SenC 86.8 74.3 47.0 43.6 48.4 86.0
Vocabulary Complexity VocC 74.6 71.7 63.1 47.7 62.6 65.5
Table 48: The specific experimental results of the IU task in the FRAUAs-OOD are used to evaluate generalization in UAs.
Aspect Abbr GPT-4o Claude-3.5 Qwen2VL-72B Qwen2VL-7B Ours
Clarity Cla 94.2 88.6 79.4 57.2 92.8
Coherence Coh 93.8 90.5 84.2 55.4 92.9
Fluency Flu 97.1 91.3 84.5 50.6 95.3
Grammaticality Gra 92.6 85.1 76.2 42.1 88.0
Image Repetition ImgR 82.5 71.6 39.9 38.5 84.7
Readability Rea 95.3 89.6 83.8 51.8 95.1
Scene Consistency SceC 83.9 73.4 34.8 30.8 89.4
Semantic Consistency SemC 85.9 73.8 40.0 29.6 85.8
Table 49: The specific experimental results of the ITIG task in the FRAUAs-OOD are used to evaluate generalization in UAs.
@@
D.2 Staged Training Experiment Results
During GenEval training, we compared unified training and staged training across tasks. For staged training, we group tasks with similar output modalities: NLG and IU as Text, IG as Image, and ITIG as Inter. Using the fine-tuning method from Appendix C.1, we train models in stages, swapping the order of these task types, resulting in six models. Experiments on the FRA-ID and FRA-ID-H (Figure 14 and Table 51) show that the model performs best on tasks from the last-trained category. In contrast, unified training provided more balanced performance across tasks.
Figure 14: Experimental results in the FRA-ID and FRA-ID-H using different types of training set orders for staged training.
Method GPT-4o Human